OpenStudy (anonymous):

Write the equation y = 2 - 4x in standard form

OpenStudy (anonymous):

what is standard form?

OpenStudy (anonymous):

y - 2 = 4(x - 0) 4x - y = -2 4x + y = 2 4x + 2y = 0

OpenStudy (anonymous):

^^ that's doesn't seem to be correct.

OpenStudy (anonymous):

4x+y=2